# Dose-aware Diffusion Model for 3D Low-count PET: Multi-institutional Validation with Reader Study and Real Low-dose Data

### v2

@@ Abstract
Enter about five key words or phrases in alphabetical order, separated by commas.
@@ I Introduction
Positron Emission Tomography (PET) is a functional imaging modality widely used in oncology, cardiology, and neurology studies [1, 2, 3]. Given the growing concern of radiation exposure and potentially increased cancer risks accompanied with PET scans, reducing the PET injection dose is desirable [4]. In addition, reducing patient motion artifacts and patients’ discomfort through shorter scans as well as enhancing image quality on lower-performance scanners would also be beneficial in clinical settings. However, reduced injection dose and shorter scan duration can negatively affect PET image quality, and may affect diagnostic performance such as the identification of low-contrast lesions [5]. Therefore, reconstructing high-quality images from noisy input is an important topic. In fact, making the radiation dose as low as reasonably achievable (the ALARA principle) is a commonly accepted practice in clinical settings.
@@
Many different deep learning (DL) methods were proposed for low-dose/low-count PET image reconstructions/restorations [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. In particular, with the advancement of cutting-edge generative models [18, 19], diffusion models have demonstrated strong potential for low-dose PET imaging. Concretely, building upon the DDPM framework [20], Gong et al. proposed to perform PET image denoising with MRI as prior information for improved image quality [21]. Unfortunately, PET-MR systems are not widely available and it would be time-consuming to obtain sequential and paired PET-MR images in reality for network training and image reconstructions. Jiang et al.[22] adopted the latent diffusion model [23] for unsupervised PET denoising. Diffusion models have also been proposed for other imaging modalities, such as CT [24] and MRI [25, 26]. However, we observe that most of these methods focus on 2D imaging, leaving the application of diffusion models for 3D PET imaging largely underexplored in the literature. Several challenges should be addressed to effectively apply diffusion models to 3D PET imaging.
First, PET imaging (PET) is intrinsically a 3D imaging modality, while previous 2D diffusion works fail to consider 3D spatial correlation between different slices. Thus, a 3D diffusion model is desirable. However, due to the hardware memory limit, directly extending the diffusion model to 3D would be difficult. There are a few previous works that aim to address the 3D imaging problems of diffusion models. For example, Chung et al.[27] proposed to apply a TV penalty term along the z-axis to remove inconsistencies within each reverse sampling step in the diffusion model. On the other hand, Lee et al.[28] utilized 2 pre-trained perpendicular 2D diffusion models to remove inconsistencies between slices. Nonetheless, presented later in this paper, these previous methods failed to produce satisfactory results for 3D low-count PET (e.g., inaccurate organ boundaries in extreme low-dose settings, inconsistent 3D reconstructions, distorted image features in some cases).
@@
Second, another challenge with PET image denoising is the high variation of image noise levels among patients, especially when dealing with multi-center, cross-scanner dataset. The noise level in PET images can be affected by various factors, including but not limited to: (1) Variations of post tracer injection acquisition start time. (2) Occasional tracer injection infiltration in some patients, causing the tracer to accumulate at the injection location, resulting in high image noise in the body. (3) The variation of patient’s BMI and associated weight-based tracer injection protocols. (4) Different hospitals using scanners with varying sensitivities have different total scan times for patients. (5) Variations of excretion into the urinary bladder. Higher excretion would also result in higher noise in other parts of the image. All these factors result in variability of noise levels in the reconstructed images. Therefore, a method to adaptively denoise images with varying noise levels is desirable. However, previously proposed methods mentioned above have limited generalizability to different noise levels. A network trained on one noise level often fails to produce high-quality reconstructions on other noise levels [16]. To address this problem, we previously proposed to combine multiple U-net-based [29] sub-networks with varying denoising power to generate optimal results for any input noise levels [16]. However, training multiple sub-networks requires tedious data pre-processing and long training time. The testing time also linearly increases with the number of sub-networks.
@@
Third, different from other imaging modalities, PET was developed as a quantitative tool, being recognized as providing an objective, and more accurate measure for prognosis and response monitoring purposes than visual inspection alone [30]. However, our experimental results showed that, although standard diffusion models (DDPM [20], DDIM[31]) produce visually appealing images, they typically fail to maintain accurate quantification and produced images with biased tracer uptake. Specifically, the tracer distribution in the entire image change after diffusion models. This issue is more severe in whole-body PET scans since the tracer uptake could vary significantly among organs.
@@
In this work, we developed a dose-aware diffusion model for 3D PET Imaging (DDPET-3D) to address these limitations. The main methodological contributions of the proposed DDPET-3D framework are: First: We proposed a 2.5D diffusion strategy with multiple fixed noise variables to address the 3D issue in diffusion models for medical imaging. DDPET-3D maintains a similar memory burden to 2D diffusion models while achieving high-quality reconstructions. Second: We proposed a dose-embedding strategy that allows noise-aware denoising. DDPET-3D can simultaneously denoise 3D PET images with varying low-dose/count levels. Third: To achieve a quantitative-accurate reconstruction, we proposed to use a denoised prior in DDPET-3D. The proposed denoised prior also allows the diffusion model to converge within 25 sampling steps. DDPET-3D can denoise 3D PET images within a reasonable time constraint ($\sim$15 mins on a single NVIDIA A40 GPU). Previous diffusion methods using DDPM sampling would require approximately $6$ hrs.
@@
Furthermore, existing models are often tailored to specific hospitals, scanners, patient populations, acquisition protocols, and noise levels, limiting their generalizability across the diverse PET image acquisition settings encountered in practice. To address this, we evaluated the proposed method on a large-scale, cross-center, cross-scanner dataset, which is potentially the most diverse and extensive dataset in PET-related AI literature. In this work, we coordinated different hospitals around the world, prepared this large-scale dataset for the evaluation of PET denoising algorithms, and have made parts of the data publicly available. In addition to traditional image quality metrics (SSIM, PSNR, etc), we designed a reader study with nuclear medicine physicians to provide a clinical insight of the model performance. This evaluation highlights the method’s clinical potential and generalizability across various acquisition settings, making it more applicable to real-world scenarios.
@@
The ultimate goal of DL methods for PET image denoising is to achieve low-dose PET without compromising image quality and clinical performance. As outlined in Fig. 1(c), evaluated on real-world, large-scale, cross-center, ultra-low-dose 3D PET data, with confirmation from reader studies with nuclear medicine physicians, DDPET-3D demonstrates superior quantitative and qualitative results that are comparable or even better than the 100% full-count data as well as previous baseline methods across varying low-count-levels. DDPET-3D also showed superior generalizability when directly applied to data acquired at a different hospital without further network fine-tuning. We also evaluated the proposed method on a cohort of human studies with real low-dose injection. Presented results show the potential of DDPET-3D to achieve low-dose PET imaging while maintaining image quality. An overview of the dataset is provided in Fig. 1(b).
@@
II Methods
II-A Diffusion models
The general idea of diffusion models is to learn the target data distribution $q({\bm{x}}_{0})$ (i.e., full-dose PET images in our case) using neural networks. Once the distribution is learned, we can synthesize a new sample from it. Diffusion models consist of two Markov chains: the forward diffusion process and the learned reverse diffusion process. The forward diffusion process $q$ gradually adds small amount of Gaussian noise to ${\bm{x}}_{0}\sim q({\bm{x}}_{0})$ in each step, until the original image signal is completely destroyed. As defined in [20]
@@ II-A Diffusion models
II-B Conditional PET image denoising
The framework described above only allows unconditional sampling. For the purpose of low-dose PET denoising, instead of generating new samples, the network needs to denoise the images based on input noisy counterparts. This can be achieved by adding additional condition to the neural network. $\epsilon_{\bm{\theta}}{({\bm{x}}_{t},t)}$ becomes $\epsilon_{\bm{\theta}}{({\bm{x}}_{t},t,{\bm{x}}_{\mathrm{noisy}})}$, where ${\bm{x}}_{\mathrm{noisy}}$ denotes the noisy input PET images. Specifically, the input to the neural network becomes a 2-channel input, one is ${\bm{x}}_{t}$, the other one is ${\bm{x}}_{\mathrm{noisy}}$. However, we noticed that, this technique results in severe inconsistencies in the reconstructed 3D image volumes. One example is presented in Fig. 9 (DDIM results).
II-C Proposed DDPET-3D
@@
The proposed framework is depicted in Fig. 2. The proposed network can observe 3D information from adjacent slices during the training process. This is achieved by using the $n$ neighboring slices as input to predict the central slice. Ablation studies showed that the performance converges at $n=31$, which was used in this work. One may use 3D convolutional layers to replace the 2D convolutional layers in the diffusion model for an enlarged receptive field and to allow the network to observe 3D structural information. However, we found that it significantly increases memory burden and makes the network difficult to optimize. To alleviate the memory burden and allow faster convergence, we embed neighboring slices in the channel dimension. Specifically, when using only one 2D slice as conditional information ($n=1$), the input dimension is $N_{b}\times W\times W\times 2$, where the last dimension is the channel dimension, $N_{b}$ is input batch size, and $W$ is the width of the images. One channel is conditional 2D slice, and the other one is ${\bm{x}}_{t-1}$. When $n=31$, the input dimension becomes $N_{b}\times W\times W\times 32$.
@@
Such technique allows the network to observe neighboring slices for 3D image reconstruction with only incremental increase in memory burden. Also, ablation studies showed that such technique also noticeably improves the inconsistency problem for 3D imaging. This is consistent with our expectation. With $n=1$, when the network tries to predict the next slice, we observe inconsistent reconstructions because both the conditional information and the starting Gaussian noise $p({\bm{x}}_{T})$ change. When $n=31$, the starting Gaussian noise $p({\bm{x}}_{T})$ still changes, but all the 31-channel conditional information only shifts a little. The network should produce more consistent output with only subtle changes in the input.
@@
Another reason causing the inconsistency issue is because the starting Gaussian noise of the reverse process is different for different slices. The diffusion sampling strategy starts from a random location in the high-dimensional space (random Gaussian noise), and approximates the data distribution $q({\bm{x}}_{0})$ based on the trained neural network, or the score function as described in [34], through many iterations. The denoising problem is ill-posed, and we could generate an infinite number of different denoised images given the same low-count input with different starting Gaussian noise. This is beneficial for generative models to produce a wide variety of different images. However, the stochastic nature of the diffusion model would be problematic for medical image reconstruction problems because we expect neighboring slices to be consistent with each other in the 3D image volume. To address this problem, we proposed to fix the starting Gaussian noise when reconstructing all the slices in the entire 3D volume. Specifically, the starting Gaussian noise at the last time step $T$, ${\bm{x}}_{T}$ is fixed for all the slices during the sampling process, only the conditional low-count PET images change for different slices.
@@
This approach produced more consistent reconstructions along the z-axis. However, we noticed that by fixing $p({\bm{x}}_{T})$, noise-dependent artifacts would propagate to all the slices along the z-axis. One example is presented in the first sub-figure in Fig. 15 (single $\epsilon$ image). To address this issue, as presented in Fig. 2, we proposed to initialize 2 different noise variables ${\bm{\epsilon}}_{0}^{a}$ and ${\bm{\epsilon}}_{0}^{b}$, and we have 2 different ${\bm{x}}_{t-1}$ in the reverse process. Fortunately, having 2 different noise variables does not significantly increase sampling time since we only need to average them at the first reverse step. This technique effectively addresses this issue.
@@
As mentioned in the introduction section, despite the visually appealing results, the diffusion model usually produced images with inaccurate quantification. We noticed that, even though U-net-based methods failed to produce satisfactory results, especially for extremely low-dose settings, they typically maintained overall image quantification much better than diffusion models. To take advantage of that, instead of starting from random Gaussian noise, we first generate a denoised prior using a pre-trained U-net-based network, and then add Gaussian noise to it based on Equation 2. This added-noise denoised prior were used as the starting point of the reverse process. We adapted the Unified Noise-aware Network (UNN) proposed by Xie et al. [16] to generate the denoised priors.
@@
For injected dose embedding, the values of the administered dose were first converted to Becquerel (Bq). Both $\sin$ and $\cos$ functions were then used to encode the injected dose and add with the diffusion time-steps (i.e., $t+\sin(\mathrm{dose})+\cos(\mathrm{dose})$, where $t$ is the diffusion time-step). Time-step $t$ was also encoded using $\sin$ and $\cos$ functions. The encoded values were then fed into 2 linear layers to generate the embedding with Sigmoid linear unit (SiLU) in between.
@@
In summary, compared to previous methods, DDPET-3D has the following contributions: (1) 3D reconstructions with fine details and addresses the 3D inconsistency issue in previous diffusion models. This was achieved by using multiple neighboring slices as conditional information and enforcing the same Gaussian latent for all slices in sampling. (2) DDPET-3D maintains accurate image quantification by using a pre-trained denoised prior in sampling. (3) DDPET-3D achieves dose-aware denoising. It can be generalized to different low-count levels. (4) With all the proposed strategies, DDPET-3D converges within 25 diffusion steps, allowing fast reconstructions. It takes roughly 15 minutes to reconstruct the entire 3D volume, while DDPM sampling takes about 6 hours.
@@
II-D PET data acquisitions and reconstructions
The collection of the patient datasets in each medical center was approved by the institutional review board (IRB) at each institution. All data in this study were de-identified prior to model training, testing and reader studies.

### v3

# Dose-aware Diffusion Model for 3D PET Image Denoising: Multi-institutional Validation with Reader Study and Real Low-dose Data

@@ Abstract
Diffusion models, PET denoising, Low-dose imaging
@@ I Introduction
Positron Emission Tomography (PET) is a functional imaging modality widely used in oncology, cardiology, and neurology studies [1, 2, 3]. Given the growing concern of radiation exposure and potentially increased cancer risks associated with PET scans, reducing the PET injection dose is desirable [4]. In addition, reducing patient motion artifacts and patient discomfort through shorter scans as well as enhancing image quality on lower-performance scanners would also be beneficial in clinical settings. However, reduced injection dose and shorter scan duration can negatively affect PET image quality, and may affect diagnostic performance such as the identification of low-contrast lesions [5]. Therefore, reconstructing high-quality images from noisy input is an important topic. In fact, making the radiation dose as low as reasonably achievable (the ALARA principle) is a commonly accepted practice in clinical settings.
@@
Many different deep learning (DL) methods were proposed for low-dose/low-count PET image reconstructions/restorations [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. In particular, with the advancement of cutting-edge generative models [18, 19], diffusion models have demonstrated strong potential for low-dose PET imaging [20, 21]. Diffusion models have also been proposed for other imaging modalities, such as CT [22] and MRI [23, 24]. However, we observe that most of these methods focus on 2D imaging, leaving the application of diffusion models for 3D PET imaging largely underexplored in the literature. Several challenges should be addressed to effectively apply diffusion models to 3D PET imaging.
First, PET imaging (PET) is intrinsically a 3D imaging modality, whereas previous 2D diffusion works do not take into account the 3D spatial correlation between different slices. Thus, a 3D diffusion model is desirable. However, due to the hardware memory limit, directly extending the diffusion model to 3D would be difficult.
@@
Second, another challenge with PET image denoising is the high variation of image noise levels among patients, especially when dealing with multi-center, cross-scanner dataset. The noise level in PET images can be affected by various factors, including, but not limited to: (1) Variations of post tracer injection acquisition start time. (2) Occasional tracer injection infiltration in some patients, which causes the tracer to accumulate at the injection location, resulting in high image noise in the body. (3) The variation of patient’s BMI and associated weight-based tracer injection protocols. (4) Different hospitals using scanners with varying sensitivities have different total scan times for patients. (5) Variations of excretion into the urinary bladder. Higher excretion would also result in higher noise in other parts of the image. All of these factors result in variability of the noise levels in the reconstructed images. Therefore, a method to adaptively denoise images with varying noise levels is desirable. However, previously proposed methods mentioned above have limited generalizability to different noise levels. A network trained on one noise level often fails to produce high-quality reconstructions on other noise levels.
@@
Third, different from other imaging modalities, PET was developed as a quantitative tool, being recognized as providing an objective, and more accurate measure for prognosis and response monitoring purposes than visual inspection alone [25]. However, our experimental results showed that although standard diffusion models (DDPM [26], DDIM[27]) produce visually appealing images, they typically fail to maintain accurate quantification and produced images with biased tracer uptake. Specifically, the tracer distribution in the entire image changes after diffusion models. This issue is more severe in whole-body PET scans since the tracer uptake could vary significantly among organs.
@@
In this work, we develop a dose-aware diffusion model for 3D PET Imaging (DDPET-3D) to address these limitations. The main methodological contributions of the proposed DDPET-3D framework are: First: We proposed a 2.5D diffusion strategy with multiple fixed noise variables to address the 3D issue in diffusion models for medical imaging. DDPET-3D maintains a similar memory burden to 2D diffusion models while achieving high-quality reconstructions. Second: We proposed a dose-embedding strategy that allows noise-aware denoising. DDPET-3D can simultaneously denoise 3D PET images with varying low-dose/count levels. Third: To achieve a quantitative-accurate reconstruction, we proposed to use a denoised prior in DDPET-3D. The proposed denoised prior also allows the diffusion model to converge within 25 sampling steps. DDPET-3D can denoise 3D PET images within a reasonable time constraint ($\sim$15 mins on a single NVIDIA A40 GPU). Previous diffusion methods using DDPM sampling would require approximately $6$ hr.
@@
Furthermore, existing models are often tailored to specific hospitals, scanners, patient populations, acquisition protocols, and noise levels, limiting their generalizability across the diverse PET image acquisition settings encountered in practice. To address this, we evaluated the proposed method on a large-scale, cross-center, cross-scanner dataset, which is potentially the most diverse and extensive dataset in PET-related AI literature. In this work, we coordinated different hospitals around the world, prepared this large-scale dataset for the evaluation of PET denoising algorithms, and have made parts of the data publicly available. In addition to traditional image quality metrics (SSIM, PSNR, etc.), we designed a reader study with nuclear medicine physicians to provide a clinical insight of the model performance. This evaluation highlights the method’s clinical potential and generalizability across various acquisition settings, making it more applicable to real-world scenarios.
@@
The ultimate goal of DL methods for PET image denoising is to achieve low-dose PET without compromising image quality and clinical performance. As outlined in Fig. 1(c), evaluated on real-world, large-scale, cross-center, ultra-low-dose 3D PET data, with confirmation from reader studies with nuclear medicine physicians, DDPET-3D demonstrates superior quantitative and qualitative results that are comparable or even better than the 100% full-count data as well as previous baseline methods across varying low-count-levels. DDPET-3D also showed superior generalizability when directly applied to data acquired at a different hospital without further network fine-tuning. We also evaluated the proposed method on a cohort of human studies with real low-dose injection. The results presented show the potential of DDPET-3D to achieve low-dose PET imaging while maintaining image quality. An overview of the dataset is provided in Fig. 1(b).
@@
II Related Work
Deep learning methods for PET image denoising can be classified into three broad categories: (1) convolutional-based models, (2) transformer-based models, and (3) diffusion-based models.
Convolutional-based models were among the first approaches to low-dose PET denoising. Many supervised networks were proposed to learn a mapping between noisy low-quality and high-quality images, including U-net variants [6, 13, 17, 28, 29], residual learning [14], and generative adversarial networks (GANs) [8, 9, 30]. While these methods show promising visual performance, they typically fail to generalize across different acquisition settings (e.g., different low-dose levels). To address this, we previously proposed a Unified Noise-Aware Network (UNN) [16] to integrate multiple U-net [31] sub-networks with varying denoising power to achieve adaptive and noise-aware denoising. However, training multiple sub-networks requires tedious data preprocessing and a long training time. The testing time also linearly increases with the number of sub-networks. Unsupervised and self-supervised methods have also been explored in the literature [32, 33, 34, 15], but they are beyond the scope of this paper.
Transformer-based models have been explored in PET image denoising mainly due to their ability to model long-range dependencies and capture global contextual information. For example, Jang et al. presented Spach Transformer [35], an efficient spatial and channel-wise encoder decoder transformer that can leverage spatial and channel information based on local and global global multi-head self-attention. Yang et al. introduced AMIR (All-In-One Medical Image Restoration) network [36], a transformer-based architecture to address different medical imaging problems using one unified model.
Diffusion-based models have demonstrated strong potential for PET image denoising by progressively refining noisy inputs through iterative denoising steps rooted in probabilistic modeling. For example, based on the DDPM framework [26], Gong et al. proposed to perform PET image denoising with MRI as prior information to improve image quality [20]. Unfortunately, PET-MR systems are not widely available and it would be time consuming to obtain sequential and paired PET-MR images in reality for network training and image reconstructions. Jiang et al.[21] adopted the latent diffusion model [37] for unsupervised PET denoising.
As mentioned above, directly training a 3D diffusion model for 3D imaging is difficult due to hardware limit. Several previous works aim to address the 3D imaging problems of diffusion models. For example, Chung et al.[38] proposed to apply a TV penalty term along the z-axis to remove inconsistencies within each reverse sampling step in the diffusion model. On the other hand, Lee et al.[39] used two pretrained perpendicular 2D diffusion models to remove inconsistencies between slices. However, presented later in this paper, these previous methods did not produce satisfactory results for 3D low-count PET (e.g., inaccurate organ boundaries in extreme low-dose settings, inconsistent 3D reconstructions, and distorted image features in some cases). Yoon et al. presented Conditional Score-based Residual Diffusion (CSRD) [40] which adapted a patch-based training strategy to alleviate the memory bottleneck associated with the 3D diffusion model. While CSRD provides a feasible path toward a volumetric diffusion model, due to the high memory demands of pure 3D diffusion models, the largest path size we could process into the GPU was smaller than the 2.5D conditional volume used in the proposed DDPET-3D method, potentially limiting its effectiveness on whole-body PET data.
In this paper, representative methods in each category were selected and compared.
III Methods
III-A Preliminary: Diffusion models
The general idea of diffusion models is to learn the target data distribution $q({\bm{x}}_{0})$ (i.e., full-dose PET images in our case) using neural networks. Once the distribution is learned, we can synthesize a new sample from it. Diffusion models consist of two Markov chains: the forward diffusion process and the learned reverse diffusion process. The forward diffusion process $q$ gradually adds a small amount of Gaussian noise to ${\bm{x}}_{0}\sim q({\bm{x}}_{0})$ at each step, until the original image signal is completely destroyed. As defined in [26]
@@ III-A Preliminary: Diffusion models
III-B Conditional PET image denoising
The framework described above only allows unconditional sampling. For the purpose of low-dose PET denoising, instead of generating new samples, the network needs to denoise the images based on input noisy counterparts. This can be achieved by adding an additional condition to the neural network. $\epsilon_{\bm{\theta}}{({\bm{x}}_{t},t)}$ becomes $\epsilon_{\bm{\theta}}{({\bm{x}}_{t},t,{\bm{x}}_{\mathrm{noisy}})}$, where ${\bm{x}}_{\mathrm{noisy}}$ denotes the noisy input PET images. Specifically, the input to the neural network becomes a two-channel input, one is ${\bm{x}}_{t}$, the other one is ${\bm{x}}_{\mathrm{noisy}}$. However, we noticed that this technique results in severe inconsistencies in the reconstructed 3D image volumes. One example is presented in Fig. 7 (DDIM results).
III-C Proposed DDPET-3D
@@
The proposed framework is depicted in Fig. 2. The proposed network can observe 3D information from adjacent slices during the training process. This is achieved by using the $n$ neighboring slices as input to predict the central slice. Ablation studies showed that the performance converges at $n=31$, which was used in this work. One may use 3D convolutional layers to replace the 2D convolutional layers in the diffusion model for an enlarged receptive field and to allow the network to observe 3D structural information. However, we found that it significantly increases the memory burden and makes the network difficult to optimize. To alleviate the memory burden and allow faster convergence, we embed neighboring slices in the channel dimension. Specifically, when using only one 2D slice as conditional information ($n=1$), the input dimension is $N_{b}\times W\times W\times 2$, where the last dimension is the channel dimension, $N_{b}$ is the input batch size, and $W$ is the width of the images. One channel is a conditional 2D slice, and the other one is ${\bm{x}}_{t-1}$. When $n=31$, the input dimension becomes $N_{b}\times W\times W\times 32$.
@@
This technique allows the network to observe neighboring slices for 3D image reconstruction with only an incremental increase in memory burden. In addition, ablation studies showed that this technique also noticeably improves the inconsistency problem for 3D imaging. This is consistent with our expectation. With $n=1$, when the network tries to predict the next slice, we observe inconsistent reconstructions because both the conditional information and the starting Gaussian noise $p({\bm{x}}_{T})$ change. When $n=31$, the starting Gaussian noise $p({\bm{x}}_{T})$ still changes, but all the 31-channel conditional information only shifts a little. The network should produce a more consistent output with only subtle changes in the input.
@@
We hypothesized that another reason for the inconsistency issue is that the starting Gaussian noise of the reverse process is different for different slices. The diffusion sampling strategy starts from a random location in the high-dimensional space (random Gaussian noise), and approximates the data distribution $q({\bm{x}}_{0})$ based on the trained neural network, or the score function as described in [43], through many iterations. The denoising problem is ill-posed, and we could generate an infinite number of different denoised images given the same low-count input with different starting Gaussian noise. This is beneficial for generative models to produce a wide variety of different images. However, the stochastic nature of the diffusion model would be problematic for medical image reconstruction problems because we expect neighboring slices to be consistent with each other in the 3D image volume. To address this problem, we proposed to fix the starting Gaussian noise when reconstructing all the slices in the entire 3D volume. Specifically, the starting Gaussian noise at the last time step $T$, ${\bm{x}}_{T}$ is fixed for all the slices during the sampling process, only the conditional low-count PET images change for different slices.
@@
This approach produced more consistent reconstructions along the z-axis (as demonstrated in Fig. 16). However, we noticed that by fixing $p({\bm{x}}_{T})$, there are some noise-dependent artifacts that propagate to all the slices along the z-axis. An example is presented in the first sub-figure in Fig. 16 (single $\epsilon$ image). To address this issue, as presented in Fig. 2, we proposed to initialize 2 different noise variables ${\bm{\epsilon}}_{0}^{a}$ and ${\bm{\epsilon}}_{0}^{b}$, and we have 2 different ${\bm{x}}_{t-1}$ in the reverse process. It is worth noting that having 2 different noise variables does not significantly increase sampling time since we only need to average them at the first reverse step. This technique effectively addresses this issue.
@@
As mentioned in the Introduction, despite the visually appealing results, the diffusion model usually produced images with inaccurate quantification. We noticed that, even though U-net-based methods failed to produce satisfactory results, especially for extremely low-dose settings, they typically maintained overall image quantification much better than diffusion models. To take advantage of that, instead of starting from random Gaussian noise, we first generate a denoised prior using a pre-trained U-Net-based network. Gaussian noise is then added to this prior following Equation 2, and the resulting perturbed image serves as the starting point for the reverse diffusion process. We adapted the Unified Noise-aware Network (UNN) proposed by Xie et al. [16] to generate the denoised priors.
Intuitively, one can imagine that diffusion models aim to approximate a very narrow distribution centered around real PET images ($q({\bm{x}}_{0})$). Starting from a high-dimensional Gaussian space makes it difficult for the model to accurately converge to this narrow distribution. In contrast, U-Net outputs provide a strong initial guess that lies closer to the true PET distribution, thereby significantly improving both visual fidelity and quantitative accuracy. This benefit is clearly demonstrated in Fig. 13 and Table IV.
@@
For the embedding of the injected dose, the values of the administered dose were first converted to Becquerel (Bq). The functions $\sin$ and $\cos$ functions were then used to encode the injected dose and add to the diffusion time steps (i.e., $t+\sin(\mathrm{dose})+\cos(\mathrm{dose})$, where $t$ is the diffusion time-step). The time step $t$ was also encoded using $\sin$ and $\cos$ functions. The encoded values were then fed into two linear layers to generate the embedding with a Sigmoid linear unit (SiLU) in between.
@@
In summary, compared to previous methods, DDPET-3D has the following contributions: (1) 3D reconstructions with fine details and addresses the issue of 3D spatial inconsistency in previous diffusion models. This was achieved by using multiple neighboring slices as conditional information, enforcing the same Gaussian latent for all slices in sampling, and using multiple starting Gaussian noise. (2) DDPET-3D preserves quantitative accuracy by using a pretrained denoised prior in sampling. (3) DDPET-3D achieves dose-aware denoising. It can be generalized to different low-count levels. (4) With all the proposed strategies, DDPET-3D converges within 25 diffusion steps, allowing fast reconstructions. It takes roughly 15 minutes to reconstruct the entire 3D volume, while DDPM sampling takes about 6 hours.
@@
III-D PET data acquisitions and reconstructions
The collection of the patient datasets in each medical center was approved by the Institutional Review Board (IRB) of each institution. All data in this study were de-identified prior to model training, testing, and reader studies.